OpenStudy (steve816):

If 10 - x is the circumference, what is the equation of the area?

OpenStudy (mathmath333):

2*pi*r=10-x pi*r^2=area

OpenStudy (steve816):

So how would I do this?

OpenStudy (mathmath333):

substtitute r=(10-x)/(2*pi)

OpenStudy (steve816):

ok thanks